# IRGS: Inter-Reflective Gaussian Splatting with 2D Gaussian Ray Tracing

### v1

@@ 2 Related work
focuses on recovering geometry, material, and lighting properties from captured images, presenting a complex challenge due to high-dimensional light-surface interactions. Inspired by the success of neural radiance fields, many methods [27, 5, 35, 38, 6, 1, 19, 14, 32, 41] leverage flexible ray marching techniques and neural implicit fields for inverse rendering. Recent approaches [10, 18, 13, 26, 30, 11, 9] have adapted 3D Gaussian splatting (3DGS) to inverse rendering by assigning material-related properties to each Gaussian primitive. Leveraging the rapid rendering capabilities of 3DGS, these methods achieve much higher efficiency compared to NeRF-based techniques. However, due to the lack of an efficient Gaussian ray tracer, existing 3DGS-based methods either apply a simplified version of the rendering equation [18] or use learnable parameters to model incident light [10], which limits the accuracy of material and lighting estimation. For example, GS-IR [18] employs split-sum approximation to simplify the rendering equation and uses baked volumes to store occlusion and indirect illumination. R3DG [10] shades each Gaussian individually, using point-based ray tracing to bake occlusion and spherical harmonics to parameterize indirect radiance. To overcome these limitations, we propose the IRGS framework that utilizes 2D Gaussian ray tracing to accurately model inter-reflections.
@@ 3.3.1 Stage I: 2D Gaussian pretraining
Training loss
Following 2DGS [12], we apply a normal consistency loss $\mathcal{L}_{\mathrm{n}}$ to ensure alignment between the rendered normal map and the underlying geometry, as well as a depth distortion loss $\mathcal{L}_{\mathrm{d}}$ to encourage concentration of the 2D Gaussians. Additionally, we introduce an edge-aware smoothness loss $\mathcal{L}_{\mathrm{s,n}}$ on the normal map to enhance geometric robustness:
@@
where $\mathcal{C}_{\mathrm{gt}}$ is the ground truth training image. To address the presence of floaters during standard 2DGS training, we also incorporate a binary cross-entropy loss to constrain the geometry using the provided object mask $\mathcal{M}$:
@@ 3.3.2 Stage II: inverse rendering
Rasterization
As introduced in Section 3.1, a dielectric BRDF model is used for the rendering equation. We assign each 2D Gaussian primitive additional material properties: diffuse albedo $\bm{a}$ and roughness $r$. Pixel-level albedo map $\mathcal{\bm{A}}$ and roughness map $\mathcal{R}$ are obtained through alpha-blending during rasterization: $\{\mathcal{\bm{A}},\mathcal{R}\}=\sum_{i=1}^{N}w_{i}\{\bm{a}_{i},r_{i}\}$.
@@
Light parametrization
We decompose the incident radiance at surface point $\bm{x}$ with direction $\bm{\omega}_{i}$ in Eq. 3 into direct radiance from infinity and indirect radiance:
@@
Rendering
From the rendered depth and normal maps, we can determine the surface position $\bm{x}$ and the corresponding surface normal vector $\bm{n}$ for each pixel coordinate. Monte Carlo sampling is used to evaluate the rendering equation. Since we focus on reconstructing objects with dielectric materials, we employ stratified sampling to uniformly sample $N_{\mathrm{r}}$ incident directions across the hemisphere. The final rendering result is then computed as:
@@
Training loss
In addition to the training losses from the first stage, we apply regularization to facilitate material and lighting estimation. Given the prior that diffuse light $L_{\mathrm{diffuse}}=\frac{1}{N_{\mathrm{r}}}\sum_{i=1}^{N_{\mathrm{r}}}L(\bm{\omega}_{i},\bm{x})$ exhibits a natural white radiance, we minimize the difference between the RGB values [19]:
@@
Relighting
When performing relighting, changes in environmental lighting make the optimized radiance $\bm{c}$ of each Gaussian no longer applicable for indirect radiance in 2D Gaussian ray tracing. Unlike previous methods [10, 18] that omit indirect lighting during relighting, we propose using the split-sum approximation [23] to efficiently query indirect radiance without recursive evaluation. Specifically, we apply ray tracing on each incident ray to alpha-blend the albedo, roughness, and normal values, resulting in $\mathcal{\bm{A}}_{\mathrm{r}}$, $\mathcal{R}_{\mathrm{r}}$, and $\mathcal{\bm{N}}_{\mathrm{r}}$. These aggregated values are then shaded using a pre-filtered environment map, following the approach in Nvdiffrec [23]. Additionally, we apply importance sampling on the intensity of the ground truth environment map instead of stratified sampling, enabling more accurate rendering equation evaluation by concentrating samples in high-intensity directions. Please refer to the appendix for further details.
@@ 5 Conclusion
Limitation
Due to the significant computational demands of rendering equation evaluation, we are unable to render at a high frame rate. However, this could be addressed by baking the incident radiance or pre-computing radiance transfer, which we leave as future work.

### v2

@@ 2 Related work
focuses on recovering geometry, material, and lighting properties from captured images, presenting a complex challenge due to high-dimensional light-surface interactions. Inspired by the success of neural radiance fields, many methods [28, 5, 36, 40, 6, 1, 20, 15, 33, 43] leverage flexible ray marching techniques and neural implicit fields for inverse rendering. Recent approaches [10, 19, 14, 27, 31, 11, 9, 37] have adapted 3D Gaussian splatting (3DGS) to inverse rendering by assigning material-related properties to each Gaussian primitive. Leveraging the rapid rendering capabilities of 3DGS, these methods achieve much higher efficiency compared to NeRF-based techniques. However, due to the lack of an efficient Gaussian ray tracer, existing 3DGS-based methods either apply a simplified version of the rendering equation [19] or use learnable parameters to model incident light [10], which limits the accuracy of material and lighting estimation. For example, GS-IR [19] employs split-sum approximation to simplify the rendering equation and uses baked volumes to store occlusion and indirect illumination. R3DG [10] shades each Gaussian individually, using point-based ray tracing to bake occlusion and spherical harmonics to parameterize indirect radiance. To overcome these limitations, we propose the IRGS framework that utilizes 2D Gaussian ray tracing to accurately model inter-reflections.
@@ 3.3.1 Stage I: 2D Gaussian pretraining
Training loss Following 2DGS [13], we apply a normal consistency loss $\mathcal{L}_{\mathrm{n}}$ to ensure alignment between the rendered normal map and the underlying geometry, as well as a depth distortion loss [13] $\mathcal{L}_{\mathrm{d}}$ to encourage concentration of the 2D Gaussians. Additionally, we introduce an edge-aware smoothness loss [10, 12] $\mathcal{L}_{\mathrm{s,n}}$ on the normal map to enhance geometric robustness:
@@
where $\mathcal{C}_{\mathrm{gt}}$ is the ground truth training image. To address the presence of floaters during standard 2DGS training, we also incorporate a binary cross-entropy loss [29] to constrain the geometry using the provided object mask $\mathcal{M}$:
@@ 3.3.2 Stage II: inverse rendering
Rasterization As introduced in Section 3.1, a dielectric BRDF model is used for the rendering equation. We assign each 2D Gaussian primitive additional material properties: diffuse albedo $\bm{a}$ and roughness $r$. Pixel-level albedo map $\mathcal{\bm{A}}$ and roughness map $\mathcal{R}$ are obtained through alpha-blending during rasterization: $\{\mathcal{\bm{A}},\mathcal{R}\}=\sum_{i=1}^{N}w_{i}\{\bm{a}_{i},r_{i}\}$.
@@
Light parametrization We decompose the incident radiance at surface point $\bm{x}$ with direction $\bm{\omega}_{i}$ in Eq. 3 into direct radiance from infinity and indirect radiance:
@@
Rendering From the rendered depth and normal maps, we can determine the surface position $\bm{x}$ and the corresponding surface normal vector $\bm{n}$ for each pixel coordinate. Monte Carlo sampling is used to evaluate the rendering equation. Since we focus on reconstructing objects with dielectric materials, we employ stratified sampling to uniformly sample $N_{\mathrm{r}}$ incident directions across the hemisphere. The final rendering result is then computed as:
@@
Training loss In addition to the training losses from the first stage, we apply regularization to facilitate material and lighting estimation. Given the prior that diffuse light $L_{\mathrm{diffuse}}=\frac{1}{N_{\mathrm{r}}}\sum_{i=1}^{N_{\mathrm{r}}}L(\bm{\omega}_{i},\bm{x})$ exhibits a natural white radiance, we minimize the difference between the RGB values [20]:
@@
Relighting When performing relighting, changes in environmental lighting make the optimized radiance $\bm{c}$ of each Gaussian no longer applicable for indirect radiance in 2D Gaussian ray tracing. Unlike previous methods [10, 19] that omit indirect lighting during relighting, we propose using the split-sum approximation [24] to efficiently query indirect radiance without recursive evaluation. Specifically, we apply ray tracing on each incident ray to alpha-blend the albedo, roughness, and normal values, resulting in $\mathcal{\bm{A}}_{\mathrm{r}}$, $\mathcal{R}_{\mathrm{r}}$, and $\mathcal{\bm{N}}_{\mathrm{r}}$. These aggregated values are then shaded using a pre-filtered environment map, following the approach in Nvdiffrec [24]. Additionally, we apply importance sampling on the intensity of the ground truth environment map instead of stratified sampling, enabling more accurate rendering equation evaluation by concentrating samples in high-intensity directions. Please refer to the appendix for further details.
@@ 5 Conclusion
Limitation Due to the significant computational demands of rendering equation evaluation, we are unable to render at a high frame rate. However, this could be addressed by baking the incident radiance or pre-computing radiance transfer, which we leave as future work.
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).